Find the real roots of the equation x^3 + x^2 - 100= 0 by the iteration method


1
Expert's answer
2021-05-31T17:32:28-0400

"4^3+4^2-100<0"

"5^3+5^2-100>0"

So, the root is between 4 and 5.


"x^2(x+1)=100"

"x=\frac{10}{\sqrt{x+1}}"

Let the initial approximation be "x_0=4.5"

"x_1=10/\sqrt{4.5+1}=4.26"

"x_2=10/\sqrt{4.26+1}=4.35"

"x_3=10/\sqrt{4.35+1}=4.32"

"x_4=10/\sqrt{4.32+1}=4.33558"

"x_5=10/\sqrt{4.33558+1}=4.32921"

"x_6=10/\sqrt{4.32921+1}=4.33180"

"x_7=10/\sqrt{4.33180+1}=4.33117"

Since the difference between x6 and x7 are very small, so the root is 4.33117